OpenStudy (lilbuddy13):

2(x-y) over square root x minus square root of y

OpenStudy (3mar):

What is the question?

OpenStudy (3mar):

\[\frac{ 2(x-y) }{ \sqrt{x}-\sqrt{y} }\]

OpenStudy (lilbuddy13):

yes that is the question. Can you please help me?

OpenStudy (3mar):

This is a forumla....WHERE IS THE QUESTION???

OpenStudy (sshayer):

\[ \frac{ 2\left( x-y \right) }{ \sqrt{x}-\sqrt{y} }=\frac{ 2\left\{ \left( \sqrt{x} \right)^2-\left( \sqrt{y} \right)^2 \right\} }{ \sqrt{x} -\sqrt{y}}\] \[=\frac{ 2\left( \sqrt{x}+\sqrt{y} \right)\left( \sqrt{x}-\sqrt{y} \right) }{ \left( \sqrt{x} -\sqrt{y}\right) }=?\]

OpenStudy (lilbuddy13):

it says to rationalize the denominator

OpenStudy (sshayer):

ok ,you can do that way also. multiply the numerator and denominator by \[\sqrt{x}+\sqrt{y}\]

OpenStudy (3mar):

so it wants simplification

OpenStudy (lilbuddy13):

okay so that would be \[2(\sqrt{x}+\sqrt{y}) \over (\sqrt{x}-\]

OpenStudy (lilbuddy13):

oops minus square root of y

OpenStudy (lilbuddy13):

Thanks sooo much!!!

OpenStudy (sshayer):

as i have done\[\sqrt{x}-\sqrt{y}\] cancels from the numerator and denominator

OpenStudy (lilbuddy13):

Okay how about this... Rationalize the numerator \[\sqrt{x}-\sqrt{x+h} \over h \sqrt{x}\sqrt{x+h}\]

OpenStudy (sshayer):

multiply the numerator and denominator by \[(\sqrt{x}+\sqrt{x+h})\]

OpenStudy (lilbuddy13):

okay thanks